s D tl K) Er D Final exam 2014-Saved to my Mac ome Insert Draw Design Layout References Mailings ReviewView ? Tell me what you want to do Share te No Spacing Heading 1 Styles Pane 12 Pis Po(1+ g)15 $34.91 (1 .04)15 $62.88 Different Dividend Growth Rates Question 4 H is growing quickly. Dividends are expected to grow at a 25 percent rate for the next three years, with the growth rate falling off to a constant 4.5 percent thereafter. If the required return is 10 percent and the company just paid a $2.40 dividend, what is the current share price? Answer 4 of 4 1of 1279 words English (United Kingdom) MacBook
Answer #1
D(t+1) = D(t)*(1+g1)
D0 2.4
For the first three years
g1 0.25
D1 3
D2 3.75
D3 4.6875
Find the price of the stock in year 4
g2 0.045
D4 = D3*(1+g2)
D4 4.898438
According to the dividend growth model.
P4= D4/(R-g2)
where R is .10
P4 89.0625
The value of the stock today = sum of present value of future cash flows.
Using R = .10
Year 1 2 3 4
Cash flow 3 3.75 4.6875 89.0625
Present value 2.73 3.10 3.52 66.9140
sum of present values 76.26
The current share price is equal to $76.26.